# SLoRD: Structural Low-Rank Descriptors for Shape Consistency in Vertebrae Segmentation

### v2

@@ 4.2 Performance Comparison
4.2.1 VerSe 2019
To evaluate the efficacy of SLoRD, quantitative experimental results are first conducted on VerSe 2019. As illustrated by Table 1, our method achieves SOTA performance on the average Dice and Hausdorff distance (HD) of the hidden test dataset compared with other single-stage, two-stage, and multi-stage methods. Specifically, our framework surpasses MedNeXt [32] (the best single-stage model) by $2.42\%\uparrow$ average Dice score and $5.38mm\downarrow$ average Hausdorff distance. And these two metrics come to $1.39\%\uparrow$ and $0.99mm\downarrow$ when compared with Spine-Transformer [38] (the strongest two-stage model), $1.53\%\uparrow$ and $1.11mm\downarrow$ compared with Payer’s method [34] (the powerful multi-stage model). Moreover, the median Dice and HD of our method also reach a quite promising performance.
4.2.2 VerSe 2020
@@
Table 3: The model efficiency evaluation between our method and other approaches on the hidden test datasets of VerSe 2019.
@@ 4.2.4 Model Efficiency
Besides, the efficiency analysis is conducted between our framework and other typical methods on the hidden test dataset of VerSe 2019. Specifically, we choose model parameters, computational costs, and inference time per case as quantitative metrics. As illustrated by Table 3, our approach only requires $25.18$M network parameters, which is the most lightweight model in the referenced methods. Besides, our approach achieves a $142.8$s average inference time for each CT scan, much faster than Payer’s method ($260.0$s) and Tao’s method ($176.3$s).
Table 4: Ablation study on the orientation of positive Z-axis, sampling interval, and the number of basic contour descriptors (ASD [44]: average surface distance).
@@
4.3 Discussion
4.3.1 Settings for the spherical-based contour representations
We investigate a detailed ablation analysis on the contour representations based on the spherical coordinate system. For a simple yet effective evaluation, we adopt the average surface distance (ASD) between contour points and the boundary to assess the contour restoration performance. The ASD value in Table 4 is calculated via the linear combination of full-rank contour descriptors from SVD. Thus, this distance value stands for the best performance that can be theoretically achieved due to vertebral non-convex shape property and deviation due to integer approximation.
@@ 4.3.1 Settings for the spherical-based contour representations
The rank of basic contour descriptors: According to Table 4, 200 basic contour descriptors can already restore the original contour with sufficient precision. That is also proven in the qualitative results of Fig.3.
@@
Table 5: Ablation study on the combination of decoders on VerSe 2020.
Method Dice score (%) $\uparrow$ $HD$ (mm) $\downarrow$
Cervical Thoracic Lumbar Mean Median Cervical Thoracic Lumbar Mean Median
Ours 90.27 93.44 90.69 92.31 95.33 6.37 4.68 6.04 5.55 3.74
w/o mask decoder 87.82 90.32 88.38 90.13 94.47 8.00 6.87 7.07 7.29 5.86
w/o coefficient decoder - - - - - - - - - -
w/o center decoder 87.41 90.50 85.96 88.64 93.80 8.13 6.14 9.11 8.09 6.75
w/o positional prompts 89.58 89.83 89.62 91.22 94.85 6.86 7.09 6.19 6.63 5.03
Table 6: Ablation study on the plug-and-play property of SLoRD when applied to other single-stage, two-stage, and multi-stage methods on vertebrae segmentation. Quantitative results on the hidden test dataset of VerSe 2019 and 2020.
@@
Methods A-Dice M-Dice A-HD M-HD
@@
4.3.2 Structural ablation on SLoRD
@@
Center decoder: Since the coupling effects between spherical centers and coefficients, removing the center decoder will diminish the segmentation performance significantly, with $3.67\%\downarrow$ average Dice score and $3.07mm\uparrow$ average HD. Besides, as portrayed in Fig.10, there exists a missing detection for a specific vertebra, resulting in sequential identification errors. That phenomenon significantly weakens segmentation results on lumber vertebrae [8].
Positional prompts: The positional priors promote the regression for spherical centroids, further boost the precision of contour generation. This detection guidance is beneficial to segment the thoracic region with adjacent vertebrae.
4.3.3 Plug-and-play
We also carried out the ablation study on the plug-and-play property of SLoRD on VerSe 2019 and 2020. Besides MedNeXt [32], the proposed SLoRD will boost segmentation results achieved by Verteformer [45], Spine-Transformer [38], Payer’s method [31] as shown in Table 6. Specifically, for Verteformer on VerSe 2020, substantial improvements have been achieved on the average Dice score and HD ($2.65\%\uparrow$, $2.16mm\downarrow$). Also, SLoRD can repair segmentation results by Payer’s approach, with $2.01\%\uparrow$ average Dice score and $0.48mm\downarrow$ average HD. Qualitative results are provided in Fig.11 to demonstrate the efficacy of SLoRD in promoting segmentation consistency when plugged into single-stage, two-stage, and multi-stage methods.
@@ 4.3.4 Comparison with PCA-based shape model
As we mentioned in Section 3.1, SVD shows the best low-rank approximation property and can mine more robust shape descriptors compared with PCA. Here a straight comparison is implemented on the hidden test dataset of VerSe 2020. As illustrated in Table 9, our method outperforms DeepSSM significantly, with a $3.92\%$ higher Dice score and a $1.73mm$ lower Hausdorff distance. Also, since different vertebrae bear various numbers of contour points, it is tricky to carry out the uniform sampling process while delineating the original contour. In contrast, our method achieves a $1.117\downarrow$ ASD value, showing more satisfactory contour reconstruction results based on the structural encoding way of contour representations.
@@ 5 Conclusion
In this paper, to address the challenge of intra-vertebrae segmentation inconsistency, we devise SLoRD by jointly imposing voxel consistency and contour precision regularizations. Specifically, we propose a structural contour representation way based on the spherical coordinate system. Besides, basic contour descriptors and shape constraints facilitate regressed contour points close to vertebral surfaces. Detailed quantitative and qualitative results on VerSe 2019 and 2020 demonstrate the superiority of our method compared with other single-stage, two-stage and multi-stage models. Moreover, SLoRD serves as a plug-and-play framework to repair the segmentation in- consistency in coarse predictions from other methods.

### v3

@@ 4.2 Performance Comparison
4.2.1 Comparison baselines
For a comprehensive comparison, we list out eight powerful single-stage baselines, five SOTA two-stage or multi-stage approaches. For single-stage baselines, Lessmann’s method [1], Xiangshang’s method [6] and Verteformer [10] are specifically devised for the vertebrae segmentation task, while 3D UNet [19], nnUNet [20], UNeXt [43], Swin UNETR [44], and MedNeXt [42] serve as general segmentation frameworks feasible for bone-family anatomical structures. Besides, several two-stage and multi-stage methods achieving SOTA performance in the VerSe Challenge are also included in our benchmark, such as Payer’s method [23], Chen’s method [6] and Spine-Transformer [7]. Also, SAM-based methods [45, 49, 50] show potential in segmenting bone-like anatomical structures. Thus, SAM-Med2D [45] is chosen as a comparative baseline in the benchmark by implementing the axial-plane segmentation. It is worth mentioning that we select a bounding box enclosing vertebrae and a positive point inside vertebrae as prompts. The positive point for each axial slice is set as the centroid of coarsely segmented mask generated from the first stage network. And the label of each segmented vertebra corresponds to the positive point’s label in coarse predictions. Thus, SAM-Med2D can be considered a two-stage pipeline.
4.2.2 Results on VerSe 2019
To evaluate the efficacy of SLoRD, quantitative experimental results are first conducted on VerSe 2019. As illustrated by Table 1, our method achieves SOTA performance on the average Dice and Hausdorff distance (HD) of the hidden test dataset compared with other single-stage, two-stage, and multi-stage methods. Specifically, our framework surpasses MedNeXt [42] (the best single-stage model) by $2.42\%\uparrow$ average Dice score and $5.38mm\downarrow$ average Hausdorff distance. And these two metrics come to $1.39\%\uparrow$ and $0.99mm\downarrow$ when compared with Spine-Transformer [7] (the strongest two-stage model), $1.53\%\uparrow$ and $1.11mm\downarrow$ compared with Payer’s method [6] (the powerful multi-stage model). As shown in Table 1, our model outperforms SAM-Med2D on hidden test datasets of VerSe 2019 with a $2.07\%$ average Dice score increase, a $2.88mm$ average HD decrease. Moreover, the median Dice and HD of our method also reach quite promising performance.
4.2.3 Results on VerSe 2020
@@
Table 3: The model efficiency evaluation between our method and other approaches on the hidden test datasets of VerSe 2019. The inference speed refers to the average processing time for one CT scan.
@@ 4.2.5 Model Efficiency
Besides, the efficiency analysis is conducted between our framework and other typical methods on the hidden test dataset of VerSe 2019. Specifically, we choose model parameters, computational costs, and inference time per case as quantitative metrics. As illustrated by Table 3, our approach only contains $25.18$M network parameters, which is the most lightweight model in the referenced methods. Besides, our approach achieves a $142.8$s average inference time for each CT scan, much faster than Payer’s method ($217.9$s) and Tao’s method ($176.3$s).
4.3 Ablation Analysis
4.3.1 Settings for the spherical-based contour representations
We investigate a detailed ablation analysis on the contour representations based on the spherical coordinate system. For a simple yet effective evaluation, we adopt the average surface distance (ASD) between contour points and the boundary to assess the contour restoration performance. The ASD value in Table 5 is calculated via the linear combination of low-rank contour descriptors from SVD. Thus, this distance value stands for the best performance that can be theoretically achieved due to vertebral non-convex shape property and deviation due to integer approximation.
Table 4: Ablation analysis of the rank number, the coordinate system, and the dimension reduction method on the segmentation performance when evaluated on VerSe 2019 and 2020. $k$ means the rank number.
Settings VerSe 2019 VerSe 2020
A-Dice $\uparrow$ A-HD $\downarrow$ A-Dice $\uparrow$ A-HD $\downarrow$
Low Rank (k=200) 91.33 5.97 92.31 5.55
Full Rank (k=2664) 91.52 5.89 92.40 5.62
Spherical System 91.33 5.97 92.31 5.55
Cartesian System 85.79 11.84 87.63 10.92
SVD (k=200) 91.33 5.97 92.31 5.55
PCA (k=200) 87.61 9.49 88.27 10.60
Table 5: Ablation study on the orientation of positive Z-axis, sampling interval, and the number of basic contour descriptors (ASD [48]: average surface distance). Bold strings stand for the final settings of SLoRD for the evaluation of segmentation performance.
@@ 4.3.1 Settings for the spherical-based contour representations
Table 6: Ablation study on the combination of decoders on VerSe 2020. $D_{coe}$: coefficient decoder, $D_{cen}$: center decoder, $D_{mask}$: mask decoder, Prompts: positional Gaussian prompts.
$D_{coe}$ $D_{cen}$ $D_{mask}$ Prompts Dice score (%) $\uparrow$ $HD$ (mm) $\downarrow$
Cervical Thoracic Lumbar Mean Cervical Thoracic Lumbar Mean
80.09 82.46 79.65 81.39 19.37 16.59 17.83 17.54
87.90 88.75 86.68 87.86 10.39 8.60 9.43 9.03
87.82 90.32 88.38 90.13 8.00 6.87 7.07 7.29
89.58 89.83 89.62 91.22 6.86 7.09 6.19 6.63
87.41 90.50 85.96 88.64 8.13 6.14 9.11 8.09
90.27 93.44 90.69 92.31 6.37 4.68 6.04 5.55
@@
The rank of basic contour descriptors: According to Table 5, $200$ basic contour descriptors can already restore the original contour with sufficient precision. That is also proven in the qualitative results of Fig.3. Considering both efficacy and efficiency, the rank number is set as $200$ instead of $500$. Moreover, we report the segmentation performance of our model with the guidance of low-rank ($k=200$) and full-rank ($k=2664$) contour descriptors. As shown in Table 4, by replacing the setting of full-rank with low-rank, the average Dice score only decreases $0.09\%$ for VerSe 2020.
@@
Spherical coordinate system vs. Cartesian coordinate system: To prove the efficacy of the spherical system in vertebral contour representations, we further conduct a comparison study on the type of coordinate systems. For a fair comparison, the number of sampled contour points for shape completion is unified as $2664$. As indicated by Table 4, the segmentation performance of SLoRD conditioned on the Cartesian coordinate system is significantly inferior to that conditioned on the spherical system. As the methodology part states, the spherical coordinate system introduces implicit angular priors corresponding to $2664$ directions. Thus, networks can better reconstruct shape representations via basic contour descriptors.
PCA vs. SVD: Since spherical-based basic descriptors are utilized to boost the final segmentation performance, a quantitative analysis on the way of extracting basic contour descriptors is carried out. Specifically, our adopted SVD strategy reveals a stronger numerical stability and accuracy than PCA. As shown in Table 4, there are $3.72\%\uparrow$ and $4.04\%\uparrow$ average Dice score for hidden test datasets of VerSe 2019 and 2020. That phenomenon demonstrates that SVD is better represent nonlinear high-dimensional vertebral data.
@@
4.3.2 Structural ablation on SLoRD
Table 7: Analysis on the plug-and-play property of SLoRD when applied to other methods on vertebrae segmentation. Quantitative results on the hidden test dataset of VerSe 2019 and 2020.
@@ 4.3.2 Structural ablation on SLoRD
Methods A-Dice $\uparrow$ M-Dice $\uparrow$ A-HD $\downarrow$ M-HD $\downarrow$
@@
3D UNet [19] 81.28 87.54 16.35 14.73
3D UNet [19] + SLoRD 85.11 87.49 14.76 11.08
nnUNet [20] 86.59 92.62 12.78 12.00
nnUNet [20] + SLoRD 88.60 90.37 9.05 8.52
@@
3D UNet [19] 83.32 88.90 10.59 9.02
3D UNet [19] + SLoRD 86.16 87.97 10.39 9.85
nnUNet [20] 86.83 91.79 9.29 7.58
nnUNet [20] + SLoRD 88.07 91.95 8.47 7.11
@@
Center decoder: Since the coupling effects between spherical centers and coefficients, adding the center decoder to the coefficient decoder will enhance the segmentation performance significantly, with $6.47\%\uparrow$ average Dice score and $8.51mm\uparrow$ average HD as shown in Table 6. Besides, as portrayed in Fig.9, there exists a missing detection for a specific vertebra, resulting in sequential identification errors. That phenomenon significantly weakens segmentation results on lumbar vertebrae [51].
@@
Positional prompts: The positional priors promote the regression for spherical centroids, further boost the precision of contour generation. This detection guidance is beneficial to segment the thoracic region with adjacent vertebrae. As revealed in Table 6, by adding positional prompts to three decoders, the segmentation performance is boosted with $1.09\%\uparrow$ Dice score and $1.08mm\downarrow$ HD.
@@
4.3.3 Analysis on the plug-and-play property
We also carried out the ablation study on the plug-and-play property of SLoRD on the hidden test dataset of VerSe 2019 and 2020. Besides MedNeXt [42], the proposed SLoRD will boost segmentation results achieved by 3D UNet [19], nnUNet [20], Verteformer [10], Swin UNETR [44], Spine-Transformer [7], Payer’s method [23] as shown in Table 7. Specifically, for Verteformer on VerSe 2020, substantial improvements have been achieved on the average Dice score and HD ($2.65\%\uparrow$, $2.16mm\downarrow$). Also, SLoRD can repair segmentation results by Payer’s approach, with $2.01\%\uparrow$ average Dice score and $0.48mm\downarrow$ average HD. These two metrics reach $0.59\%\uparrow$ and $1.02mm\downarrow$ for Tao’s method. Qualitative results are provided in Fig.10 to demonstrate the efficacy of SLoRD in promoting segmentation consistency when plugged into single-stage, two-stage, and multi-stage methods.
Figure 12: (a) The mask refinement results by SLoRD even with segmentation inconsistency in the topmost vertebra. (b) The limitations of SLoRD. Red: mask adhesion between thoracic vertebrae. Yellow: Imprecise contour regression for the articular processes.
4.4 Discussion
4.4.1 The challenge of similar appearances, various pathologies and metal artifacts
Talking about how to address the identification and segmentation of vertebrae with similar appearances, our proposed SLoRD adopts sequential modeling for three consecutive vertebrae. The introduction of Gaussian positional prompts helps to distinguish vertebrae with similar appearances. Besides, basic contour descriptors further boost the precision of contour regression. By labeling sequential binary masks, we could realize precise vertebrae segmentation with arbitrary FOVs.
For pathological vertebrae with deformed shapes, coarse predictions in the first stage may not ensure accurate geometric centers. However, as mentioned in the details of the center decoder, coarse centers will be refined as spherical centroids under the regularization of $\mathcal{L}_{center}$ as revealed in Fig.4. Additionally, the coupling between spherical centroids and linear coefficients is strongly correlated with the vertebral shape. By jointly optimizing spherical centers and contour points, SLoRD is capable to better detect geometric centers of vertebrae even with various pathologies.
Besides, for the precise segmentation of CT scans with metal artifacts, existing methods may result in segmentation inconsistency on the exact vertebrae with metal artifacts as depicted by Fig.6 and Fig.1. But the positive side is that they can depict precise binary masks, not covering regions with metal artifacts. Here according to the mechanism of devised contour representations, the spherical system chooses the maximum distance at the same angle. Thus, when there exist holes inside the specific vertebra, the contour descriptor is robust enough to represent the exterior contour information, despite metal artifact-corrupted regions also being included. Thus, by combining coarse predictions by the first-stage network and labeled results by SLoRD, we will acquire refined segmentation results with both shape precision and excellent intra-vertebrae consistency.
4.4.2 Comparison with iterative fully convolutional network
We have mentioned that our proposed pipeline is superior to existing methods due to the synchronous introduction of contour precision and explicit consistency constraints. Herein we focus on the analysis of Lessmann’s single-stage pipeline [1] which is based on the similar strategy of iteratively sliding 3D patches for parallel tasks of vertebral detection and segmentation. According to the analysis from recent works [7] and [21], Lessmann’s method heavily relies on the first detection, and the failure to find the first segment may result in the failure to precisely detect all subsequent vertebrae. In fact, the iterative framework [1] determines the location of the topmost vertebrae using a sliding window fashion with a constant step size. Hence, the choice for the step size is tricky to acquire the first patch precisely centered at the topmost vertebra.
Table 8: Comparisons with other single-stage, two-stage, and multi-stage methods on the SPIDER MR dataset. It is worth to mention that our work is aimed at the segmentation task for bone-like structure. Thus, only the shape of vertebrae anatomies are used for evaluation, intervertebral discs and spinal cord not included.
Method SPIDER MRI
A-Dice $\uparrow$ A-HD $\downarrow$ M-Dice $\uparrow$ M-HD $\downarrow$
3D UNet [19] 91.96 6.89 94.40 5.25
nnUNet [20] 92.84 4.97 95.27 4.38
Verteformer [10] 92.60 5.93 95.03 4.83
UNeXt [43] 92.51 6.27 94.93 5.46
Swin UNETR [44] 92.44 6.52 94.61 5.21
MedNeXt [42] 93.48 4.55 95.76 3.89
Payer C. [23] 93.19 4.28 95.60 3.74
Tao R. [7] 93.46 3.95 95.87 3.36
Ours 94.75 3.11 96.39 2.58
In comparison, our two-stage framework detects the topmost vertebra via the label distribution from predictions in the first stage, which can provide sufficient positional information. In the second stage, SLoRD can jointly optimize the vertebral detection and segmentation tasks. The interaction between the spherical center and contour points will refine the detection precision of vertebrae and boost the segmentation performance in the meantime. Thus, even if there is a label inconsistency in the top two vertebrae as shown in Fig.12(a), SLoRD is robust enough to output segmentation masks with good consistency.
@@ 4.4.3 Comparison with PCA-based shape model
As we mentioned in Section 3.1, SVD shows the best low-rank approximation property for high-dimensional nonlinear data and can mine more robust shape descriptors compared with PCA. Here a straight comparison is implemented on the hidden test dataset of VerSe 2020. As illustrated in Fig.11, our method outperforms DeepSSM significantly, with a $3.92\%$ higher Dice score and a $1.73mm$ lower Hausdorff distance. Also, since different vertebrae bear various numbers of contour points, it is tricky to carry out the uniform sampling process while delineating the original contour. In contrast, our method achieves a $1.117\downarrow$ ASD value, showing more satisfactory contour reconstruction results based on the structural encoding way of contour representations.
4.4.4 Generalizability on datasets with diverse modalities
To further demonstrate the generalization ability of SLoRD, we conduct experiments on diverse modalities of vertebral data. SPIDER is an MRI vertebral dataset different from VerSe datasets with the CT modality. As shown in Table 8, our proposed two-stage model outperforms Spine-Transformer [7] by $1.29\%\uparrow$ average Dice score and $0.84mm\downarrow$ average Hausdorff distance. Additionally, our model is superior to Payer’s method with $1.56\%\uparrow$ average Dice score and $1.17mm\downarrow$ average Hausdorff distance. Also, the median Dice and HD values of our model reach the best among the benchmark. Therefore, SLoRD reveals quite a promising generalization ability on vertebral data with different modalities.
4.4.5 The number of vertebrae to address in a sliding window by SLoRD
Since a sequential label information naturally exists inside the spine, three vertebrae are involved into the refinement process by SLoRD to boost the identification of vertebrae. Therefore, we conduct an ablation study on the number of vertebrae processed in a sliding window. As shown in Table 9, when three vertebrae are addressed in each sliding window, we can achieve the best segmentation performance, with $2.93\%\uparrow$ average Dice score compared with the setting of 1-vertebra. Besides, the setting of 2-vertebrae achieves comparable segmentation performance compared to that of 3-vertebrae. Indeed, the setting of 2-vertebrae processed in a sliding window also works due to the existence of sequential label priors.
Table 9: Performance analysis on the number of vertebrae processed each time in a sliding window by SLoRD when evaluated on 2020.
Settings VerSe 2020
A-Dice $\uparrow$ A-HD $\downarrow$ M-Dice $\uparrow$ M-HD $\downarrow$
3-vertebrae 92.31 5.55 95.33 3.74
2-vertebrae 92.10 5.95 95.17 3.90
1-vertebra 89.38 7.73 94.96 4.26
4.4.6 Limitations
Although the proposed SLoRD could achieve segmentation results with better intra-vertebrae consistency and contour precision. There still exists segmentation errors for the superior and inferior articular processes, especially in the region of with large curvature changes. Qualitative results are visualized in Fig.12(b). Moreover, there also might be a slight mask adhesion in the vertebral body, especially the thoracic vertebrae which show a close distance to each other. Under that circumstance, a uniform sampling degree cannot meet the demand for the shape modeling with complex shapes. Instead, we reckon that a curvature-based sampling strategy should be considered to deal with abrupt boundary changes in the future work.
@@ 5 Conclusion
In this work, to mitigate segmentation inconsistencies within vertebrae caused by the low contrast between neighboring vertebral structures and diverse pathological manifestations, we introduce SLoRD (Spherical-coordinate-based Local Regularization and Detection), a novel framework that simultaneously enforces voxel-level uniformity and boundary refinement constraints. Specifically, we develop an innovative structural contour encoding method grounded in spherical coordinate transformation. Additionally, fundamental contour descriptors coupled with anatomical shape priors guide the regressed boundary points to closely adhere to the vertebral surfaces. Comprehensive quantitative and qualitative evaluations on the VerSe 2019 and 2020 datasets confirm the superior performance of our approach over existing single-stage, two-stage, and multi-stage segmentation models. Furthermore, SLoRD functions as a versatile plug-and-play module, effectively correcting segmentation inconsistencies in coarse predictions generated by other methods. As for the limitation of SLoRD on segmenting regions with large curvatures, the future research is required by potentially adopting the curvature-based sampling strategy, which will further boost the segmentation performance.
Acknowledgment
Conflict of Interest: The authors declare that they have no conflict of interest.